# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Cameras

### v2

@@ 4.3 Graph Encoder
Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{c})\in\mathbb{R}^{(H\times W)\times K}$.
Graph convolution. After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the adjacency matrix $\mathbf{A}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors. We utilize the graph convolution $\mathscr{G}$ from Selvan et al. (2020) to further propagate information across the graph. Specifically, for a single graph convolution with its parameter $W_{g}$, the operation is defined as
$\mathbf{V_{c}^{{}^{\prime}}}=\mathscr{G}(\mathbf{V}_{c},\mathbf{A})=\sigma(\mathbf{A}\mathbf{V}_{c}^{T}W_{g}),$ (6)
where $\sigma(\cdot)$ is a non-linear activation function. Two serial graph convolutional networks (GCNs) are used to capture the relationships between nodes. The node features updated by the GCNs are given by
$\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\mathbf{A}),\mathbf{A}],$ (7)
@@
Graph re-projection. Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$. Since inverting the projection weight matrix is not feasible, we follow the approach in Li and Gupta (2018); Luo et al. (2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$. This is given by,
@@
Thus, $\mathscr{P}_{V\to R}$ thus linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters. Note that even if two pixels are assigned to the same node, they will have different features after re-projection. Therefore, the GCU may preserve the spatial details of the signal. Finally, these projection results can be integrated into an existing residual extractor network $\mathscr{R}$ introduced by Teed and Deng (2020). Final context feature $\mathbf{F}_{c}$ given by,
$\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{c}+\alpha\widetilde{\mathbf{R}}_{c}),$ (9)
@@
5 Experimental Results
@@
Figure 4: Visualization examples of different algorithms in Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3) are shown. For each problem, the red box in the top optical flow map and the yellow box in the bottom EPE error map indicate the same region of the flow field.
5.1 Experimental Details
SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU. We employ the Adam optimizer with a warm-up strategy during training. The model parameters for DIP are consistent with those specified in the original paperZheng et al. (2022). For each problem, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs. The number of iterative refinements is set to 10, and the batch size is 4.
Following Cai et al. (2019b) and Lagemann et al. (2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms. Additionally, given the absence of a dedicated outlier metric in PIV, we introduced Root Mean Squared Error (RMSE) as a supplementary metric due to its sensitivity to outliers.
Since SIV outputs a prediction sequence of length $N$ at each iteration, the loss function is defined as:
$loss=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1},w_{i}=\gamma^{N-i-1}.$ (10)
We set $\gamma=0.8$ during training.
@@ 5.1 Experimental Details
Problem 1 SIV (ours) 0.242 / 0.284 0.157 / 0.182 0.918 / 1.288 0.574 / 0.765 0.638 / 0.934 0.407 / 0.563 0.700 / 1.076 0.433 / 0.616 0.625 / 0.895 0.393 / 0.532
RAFT-PIV - Image 0.377 / 0.437 0.215 / 0.249 2.181 / 2.769 1.252 / 1.574 1.336 / 1.701 0.842 / 1.072 1.875 / 2.486 1.064 / 1.394 1.442 / 1.848 0.843 / 1.072
RAFT-PIV - Spike 0.347 / 0.401 0.208 / 0.240 1.223 / 1.570 0.755 / 0.942 0.857 / 1.136 0.557 / 0.731 0.958 / 1.288 0.581 / 0.755 0.846 / 1.099 0.525 / 0.667
PIV-NetS - Image 0.545 / 0.904 0.301 / 0.478 2.560 / 3.136 1.497 / 1.819 1.574 / 2.121 0.985 / 1.284 2.247 / 2.845 1.320 / 1.651 1.731 / 2.253 1.026 / 1.308
PIV-NetS - Spike 0.553 / 0.821 0.289 / 0.424 1.507 / 1.991 0.874 / 1.124 1.252 / 1.782 0.734 / 1.019 1.382 / 1.903 0.782 / 1.040 1.173 / 1.624 0.670 / 0.902
SCFlow 0.354 / 0.468 0.222 / 0.284 2.232 / 2.706 1.305 / 1.628 1.374 / 1.703 0.901 / 1.110 2.132 / 2.684 1.169/ 1.435 1.522 / 1.890 0.899 / 1.114
HiST-SFlow 0.369 / 0.424 0.224 / 0.256 1.271 / 1.614 0.796 / 0.980 0.964 / 1.250 0.611 / 0.784 1.029 / 1.354 0.637 / 0.809 0.908 / 1.160 0.567 / 0.708
coarse-to-fine HS 6.790 / 8.448 0.509 / 0.655 7.415 / 9.354 1.998 / 3.079 7.683 / 10.783 2.558 / 4.463 9.388 / 7.274 1.954 / 3.146 7.819 / 9.751 1.755 / 2.836
Problem 2 SIV (ours) 0.366 / 0.443 0.256 / 0.300 1.721 / 2.602 1.076 / 1.521 1.098 / 1.671 0.717 / 1.026 1.482 / 2.493 0.907 / 1.412 1.167 /1.802 0.739 / 1.065
RAFT-PIV - Image 0.665 / 0.785 0.368 / 0.434 4.730 / 5.952 2.635 / 3.297 2.320 / 2.906 1.366 / 1.707 4.305 / 5.572 2.368 / 3.065 3.005 / 3.804 1.684 / 2.126
RAFT-PIV - Spike 0.688 / 0.792 0.412 / 0.477 2.410 / 3.169 1.483 / 1.882 1.468 / 1.936 0.959 / 1.236 2.071 / 2.864 1.241 / 1.649 1.659 / 2.190 1.024 / 1.311
PIV-NetS - Image 1.163 / 1.896 0.683 / 1.029 4.866 / 5.944 2.830 / 3.409 2.673 / 3.719 1.643 / 2.207 4.347 / 5.497 2.500 / 3.095 3.262 / 4.264 1.914 / 2.435
PIV-NetS - Spike 0.887 / 1.552 0.540 / 0.836 3.172 / 4.155 1.818 / 2.315 2.392 / 3.400 1.410 / 1.963 3.067 / 4.154 1.744 / 2.316 2.379 / 3.315 1.378 / 1.857
SCFlow 0.694 / 0.716 0.403 / 0.486 4.689 / 5.975 2.613 / 3.156 2.387 / 3.055 1.391 / 1.697 4.326 / 5.262 2.380 / 3.123 3.024 / 3.752 1.696 / 2.116
HiST-SFlow 0.556 / 0.641 0.364 / 0.419 2.212 / 2.947 1.365 / 1.756 1.404 / 1.879 0.998 / 1.308 1.911 / 2.693 1.179 / 1.595 1.521 / 2.040 0.976 / 1.269
coarse-to-fine HS 20.691 / 20.993 8.244 / 9.628 17.164 / 19.627 8.051 / 9.944 17.469 / 21.594 8.170 / 11.252 17.131 / 19.572 7.936 / 9.970 18.114 / 20.447 8.100 / 10.199
Problem 3 SIV (ours) 0.241 / 0.286 0.155 / 0.181 0.903 / 1.282 0.564 / 0.761 0.621 / 0.916 0.393 / 0.542 0.698 / 1.106 0.437 / 0.634 0.616 / 0.897 0.389 / 0.530
RAFT-PIV - Image 0.447 / 0.523 0.244 / 0.286 2.694 / 3.334 1.495 / 1.847 1.447 / 1.808 0.888 / 1.104 2.311 / 2.948 1.271 / 1.613 1.725 / 2.153 0.974 / 1.213
RAFT-PIV - Spike 0.320 / 0.373 0.200 / 0.232 1.203 / 1.558 0.740 / 0.939 0.798 / 1.061 0.530 / 0.698 0.946 / 1.297 0.576 / 0.770 0.817 / 1.072 0.512 / 0.660
PIV-NetS - Image 0.667 / 0.991 0.343 / 0.513 2.786 / 3.357 1.602 / 1.924 1.641 / 2.164 0.956 / 1.258 2.393 / 2.988 1.378 / 1.709 1.872 / 2.375 1.070 / 1.351
PIV-NetS - Spike 0.476 / 0.789 0.247 / 0.406 1.420 / 1.914 0.856 / 1.134 1.157 / 1.695 0.716 / 1.021 1.342 / 1.891 0.774 / 1.063 1.099 / 1.572 0.648 / 0.906
SCFlow 0.462 / 0.743 0.269 / 0.433 1.850 / 2.206 0.837 / 1.008 1.300 / 1.915 0.774 / 1.142 1.972 / 2.618 0.769 / 1.040 1.396 / 1.871 0.662 / 0.905
HiST-SFlow 0.329 / 0.381 0.180 / 0.210 1.179 / 1.523 0.640 / 0.828 0.832 / 1.093 0.452 / 0.616 0.960 / 1.302 0.476 / 0.663 0.825 / 1.075 0.437 / 0.579
coarse-to-fine HS 7.908 / 8.946 1.044 / 1.648 7.924 / 9.634 2.706 / 3.816 8.345 / 11.002 3.130 / 4.828 7.854 / 9.668 2.716 / 3.903 8.008 / 9.813 2.399 / 3.549
[…]
5.2 Comparative Experiments
To evaluate the effectiveness of our proposed method, we conduct extensive experiments on the PSSD dataset. The experiments are performed under identical settings to ensure fairness in the results.
Index Settings Problem 1 Problem 2 Problem 3
DPHT GE $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$
(A) 0.625 0.393 1.167 0.739 0.616 0.389
(B) ✓* 0.658 0.395 1.249 0.764 0.627 0.389
(C) 0.757 0.403 1.278 0.777 0.630 0.391
(D) ✓* 0.645 0.396 1.242 0.752 0.619 0.386
(E) 0.701 0.398 1.249 0.793 0.659 0.401
(F) 0.805 0.431 1.349 0.887 0.725 0.413
Table 2: Ablation study of our proposed modules, showing results for End-Point Error (EPE). Blocks colored greener indicate better performance, characterized by lower EPE values. ✓ indicates the presence of the module, ✗ indicates the absence, and ✓* indicates the removal of the submodules under this module, for DPHT, it refers to 3D-DPD, and for GE, it refers to GCU (including the projection and re-projection). Setting (A) represents our final model.
We compare our method with multiple open-source baselines, including a traditional PIV algorithmCai et al. (2018), PIV networks (PIV-NetSCai et al. (2019b), RAFT-PIVLagemann et al. (2021)), as well as their spike versions, and spike-based optical flow networks (SCFlowHu et al. (2022), HiST-SFlowZhao et al. (2024)).
Results on Problem 1. In the steady flow scenario, the fluid moves at a moderate speed. In both the time intervals of 1 and 2, SIV achieves the smallest EPE and RMSE, providing the optimal estimation. Compared to the suboptimal methods, SIV reduces the EPE by 26.1% when $\Delta t=21$. The traditional coarse-to-fine HSCai et al. (2018) method can only yield relatively reasonable predictions when the time interval and displacement are small, which reflects, to some extent, the challenge of the dataset. Figure 4 shows the displacement field predictions (optical flow maps) and error maps for different methods in PIV. The red boxes in the optical flow map correspond to the yellow boxes in the error map. The marked region in Problem 1 is a smooth boundary between the two components. By comparing with the ground truth optical flow map, it is clear that the optical flow map shapes predicted by SIV and the spike version of RAFT-PIVLagemann et al. (2021) are mostly accurate, while other methods show varying degrees of distortion in this region. In the error map, it is more intuitive to see that the red areas (high error regions) in SIV are much fewer.
Results on Problem 2. Problem 2 represents high-speed scenarios with greater particle displacements. SIV maintains its lead in all test scenarios, with a 23.3% reduction in EPE at $\Delta t=21$ compared to the suboptimal method. The traditional method struggles to make effective predictions in this scenario. In the visualization example of Problem 2 shown in Figure 4, the marked area includes a fine flow field shaped like a ”human hand” and a ”laminar flow region” with very high flow speed. The predicted flow field shape of SIV is very close to the ground truth, and the error in the high-speed ”laminar flow region” is significantly smaller than that of other methods.
Results on Problem 3. Problem 3 reflects scenarios with uneven lighting, where areas may exist that cause overexposure in traditional cameras (this issue does not occur in spike cameras due to their principle). SIV still achieves the best performance across all sub-datasets. In the visualization example of Problem 3 (see Figure 4), only SIV predicts a flow field in the marked region that is nearly consistent with the ground truth, and has the fewest red areas (high error regions) in the error map. As we anticipated, in Problem 3, spike-based methods outperform image-based methods, reflecting the high dynamic range advantage of spike cameras under uneven lighting conditions.
5.3 Ablation Study
We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 2. Each column in the table is color-coded with a gradient from yellow to green, where the greener color represents better performance. These results underscore the effectiveness of the DPHT and Graph Encoder (GE) and their submodules. In Problem 1, both DPHT and GE exhibit notable improvements. In high-speed motion scenarios (Problem 2), particles in the fluid exhibit greater displacements between the source time and target time, increasing the difficulty of accurate flow prediction. Under these circumstances, the enhancements provided by the 3D-DPD submodule in DPHT and the GCU submodule in GE (7.0% and 6.4%, respectively) are the most significant across the three Problems. In Problem 3, the distinctive brightness distribution of the scene might aid feature extraction, and the improvements from DPHT and GE are comparatively less pronounced but still remain observable.
6 Conclusions
We start from the perspective of fluid (primarily turbulent) scenarios and conduct a detailed analysis of their prominent characteristics, which may pose challenges to PIV hardware and deep learning algorithms. Based on these analyses, we explore a novel approach to estimate fluid motion using spike cameras. Additionally, we develop a tailored network architecture called SIV, designed to solve PIV problems while adapting to scenario-specific features, which integrates two novel modules: DPHT and GE. DPHT is mainly used to aggregate motion features from the spike stream while retaining scene details. GE is primarily used to extract contextual features from highly unstructured turbulent scenes. We also propose a more challenging and realistic PIV dataset, PSSD, which features multiple scenarios and captures the temporal dynamics of flow field variations. Through comparative experiments, we validate that our method outperforms existing open-source PIV algorithms and spike-based optical flow algorithms on the PSSD dataset. In ablation studies, we gradually remove the introduced modules and their submodules to prove their effectiveness and analyze their performance in different scenarios. In summary, this study explores the great potential of spike cameras in PIV and advocates for the use of customized network solutions for fluid motion estimation tasks.

### v3

@@ Introduction
Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{s})\in\mathbb{R}^{(H\times W)\times K}$.
Graph convolution layer. After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the dynamic adjacency matrix $\widetilde{\mathbf{A}}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors. The detailed definition of GAT can be found in the supplementary material. Here, we implicitly denote the operations of the two GAT modules as $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$. The updated node features are given by
$\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\widetilde{\mathbf{A}})+\mathbf{V}_{c},\widetilde{\mathbf{A}}]+\mathscr{G}_{1}(\mathbf{V}_{c},\widetilde{\mathbf{A}})+\mathbf{V}_{c},$ (6)
@@
Graph re-projection. Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$. Since inverting the projection weight matrix is not feasible, we follow the approach in (Li and Gupta 2018; Luo et al. 2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$. This is given by,
@@
Thus, $\mathscr{P}_{V\to R}$ linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters. Note that even if two pixels are assigned to the same node, they will have different features after re-projection. Therefore, the GCU may preserve the spatial details of the signal. Finally, these projection results can be integrated into a residual extractor network $\mathscr{R}$ for local refinement. Final context feature $\mathbf{F}_{c}$ given by,
$\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{s}+\alpha\widetilde{\mathbf{R}}_{c}),$ (8)
@@
Multi-scale Iterative Optimizer
We adopt the Multi-scale Iterative Optimizer (MSIO) from DIP (Zheng et al. 2022) as the backbone for iterative residual flow estimation in SIV. Unlike RAFT’s full correlation volume construction, MSIO leverages Inverse PatchMatch to significantly reduce the computational cost introduced by large-scale feature maps (1/4 resolution of the full size). The detailed architecture of MSIO can be found in (Zheng et al. 2022); here, we denote it implicitly as $\mathscr{U}$. The output multi-scale velocity field is given by
$\{\mathbf{Vel}^{i}\}_{i=1}^{\text{iters}}=\mathscr{U}(\mathbf{F}_{c},\mathbf{F}_{m}),$ (9)
$\mathbf{F}_{m}$ denotes the motion features extracted by the Motion Encoder. iter denotes the number of iterations of GRU (Teed and Deng 2020). The detailed architecture of the Motion Encoder is provided in the supplementary material.
Multi-scale Velocity Refinement
DIP adopts the same compact and effective convex upsampling strategy as RAFT to recover full-resolution optical flow. However, such upsampling is insufficient for accurately reconstructing the velocity field of fluids. To address this, we propose a Multi-scale Velocity Refinement (MSVR) module to assist in reconstructing the complete velocity field, including fine-scale vortices. As illustrated in Figure 4, we first apply a sequence of 2D convolutions followed by activation functions to process the multi-scale velocity fields generated by MCIO. Information exchange across scales is achieved via cross convolution from coarse to fine levels. The extracted features are then fused and processed by distinct convolutional networks to generate a residual velocity field $\mathbf{Vel}^{\text{res}}$ and a quality map $\mathbf{Q}$. The refined fluid velocity field is given by,
$\mathbf{Vel}^{*}=\mathbf{Vel}^{i=\text{iters}}+\mathbf{Vel}^{\text{res}}\odot\mathbf{Q},$ (10)
$\odot$ denotes element-wise multiplication.
Figure 4: Multi-scale Velocity Refinement Module.
Loss Function
Since SIV outputs a prediction sequence of length $N$ at each iteration, the flow loss is defined as:
$L_{\text{flow}}=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1}.$ (11)
We introduce a gradient loss to prevent the generated flow field from becoming overly smooth in regions like shear layers or small vortical structures. It is computed as follows:
$L_{\text{grad}}=\sum_{i=0}^{i=N}w_{i}\cdot(\|\nabla_{\mathbf{x}}f_{i}-\nabla_{\mathbf{x}}f_{gt}\|_{1}+\|\nabla_{\mathbf{y}}f_{i}-\nabla_{\mathbf{y}}f_{gt}\|_{1}).$ (12)
We set $w_{i}=\gamma^{N-i-1}$ and $\gamma=0.8$ for both $L_{\text{flow}}$ and $L_{\text{grad}}$. The total loss can be formulated as follows:
$L=L_{\text{flow}}+\beta\cdot L_{\text{grad}}.$ (13)
Here, we set $\beta=0.3$.
Experimental Results
@@
Figure 5: Visualization examples of different algorithms in Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3) are shown. For each problem, the red box in the top optical flow map and the green box in the bottom EPE error map indicate the same region of the flow field.
Experimental Details
SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU. We employ the Adam optimizer with a warm-up strategy during training. For each sub-dataset, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs. The number of iterative refinements of GRUs is set to 12. The batch size is set to 4.
Following (Cai et al. 2019b) and (Lagemann et al. 2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms. Additionally, we incorporate another widely used PIV metric, Root Mean Squared Error (RMSE), in our extensive experiments, as detailed in the supplementary material.
@@
coarse-to-fine HS 6.790 0.509 7.415 1.998 7.683 2.558 9.388 1.954 7.819 1.755
RAFT-PIV-Image 0.377 0.215 2.181 1.252 1.336 0.842 1.875 1.064 1.442 0.843
RAFT-PIV-Spike 0.347 0.208 1.223 0.755 0.857 0.557 0.958 0.581 0.846 0.525
PIV-NetS-Image 0.545 0.301 2.560 1.497 1.574 0.985 2.247 1.320 1.731 1.026
PIV-NetS-Spike 0.553 0.289 1.507 0.874 1.252 0.734 1.382 0.782 1.173 0.670
SCFlow 0.354 0.222 2.232 1.305 1.374 0.901 2.132 1.169 1.522 0.899
HiST-SFlow 0.369 0.224 1.271 0.796 0.964 0.611 1.029 0.637 0.908 0.567
Flowformer-Spike 0.286 0.147 1.069 0.642 0.727 0.437 0.806 0.472 0.722 0.425
SIV (ours) 0.239 0.159 0.896 0.587 0.613 0.422 0.680 0.441 0.607 0.402
Table 1: EPE ($\downarrow$) results on PSSD dataset (Problem 1). The best results are marked in bold.
Comparative Experiments
We compare our method against several open-source baselines, including a classical PIV algorithm (Cai et al. 2018), PIV networks (PIV-NetS (Cai et al. 2019b), RAFT-PIV (Lagemann et al. 2021)), spike-adapted general optical flow networks (e.g., Flowformer with spike input (Huang et al. 2022)), and spike-based optical flow networks (SCFlow (Hu et al. 2022), HiST-SFlow (Zhao et al. 2024)).
Methods Channel Isotropic Mhd Mixing Average
coarse-to-fine HS 20.691 8.244 17.164 8.051 17.469 8.170 17.131 7.936 18.114 8.100
RAFT-PIV-Image 0.665 0.368 4.730 2.635 2.320 1.366 4.305 2.368 3.005 1.684
RAFT-PIV-Spike 0.688 0.412 2.410 1.483 1.468 0.959 2.071 1.241 1.659 1.024
PIV-NetS-Image 1.163 0.683 4.866 2.830 2.673 1.643 4.347 2.500 3.262 1.914
PIV-NetS-Spike 0.887 0.540 3.172 1.818 2.392 1.410 3.067 1.744 2.379 1.378
SCFlow 0.694 0.403 4.689 2.613 2.387 1.391 4.326 2.380 3.024 1.696
HiST-SFlow 0.556 0.364 2.212 1.365 1.404 0.998 1.911 1.179 1.521 0.976
Flowformer-Spike 0.533 0.284 1.975 1.179 1.214 0.748 1.668 0.972 1.348 0.796
SIV (ours) 0.370 0.254 1.687 1.080 1.071 0.722 1.449 0.912 1.144 0.742
Table 2: EPE ($\downarrow$) results on PSSD dataset (Problem 2). The best results are marked in bold.
As shown in Table 1, 2, and 3, $\Delta t=21$ and $\Delta t=11$ represent the number of timestamps for the central temporal intervals between two spike streams used for velocity field prediction, respectively. For detailed descriptions, please refer to the supplementary material. Our proposed SIV method achieves the lowest average End-Point Error (EPE) across all evaluated sub-datasets. Compared to baseline approaches, SIV demonstrates superior prediction accuracy in both low-velocity (Problem 1) and high-velocity (Problem 2) optical flow scenarios. Furthermore, in high dynamic range conditions (Problem 3), SIV effectively exploits the inherent advantages of spike-based input modalities to address the fundamental limitations of conventional frame-based imaging systems.
Figure 5 presents the visualized predictions of all methods across three datasets. The top row shows the velocity fields visualized using the standard optical flow color coding (Baker et al. 2011). The bottom row shows the endpoint error (EPE) maps between the predicted and ground-truth flow. Shear layers are often a primary source of turbulence generation and development. SIV achieves the lowest prediction error in both shear layers and small-scale vortices, while avoiding noticeable blurring, distortion, or discontinuities.
Methods Channel Isotropic Mhd Mixing Average
coarse-to-fine HS 7.908 1.044 7.924 2.706 8.345 3.130 7.854 2.716 8.008 2.399
RAFT-PIV-Image 0.447 0.244 2.694 1.495 1.447 0.888 2.311 1.271 1.725 0.974
RAFT-PIV-Spike 0.320 0.200 1.203 0.740 0.798 0.530 0.946 0.576 0.817 0.512
PIV-NetS-Image 0.667 0.343 2.786 1.602 1.641 0.956 2.393 1.378 1.872 1.070
PIV-NetS-Spike 0.476 0.247 1.420 0.856 1.157 0.716 1.342 0.774 1.099 0.648
SCFlow 0.462 0.269 1.850 0.837 1.300 0.774 1.972 0.769 1.396 0.662
HiST-SFlow 0.329 0.180 1.179 0.640 0.832 0.452 0.960 0.476 0.825 0.437
Flowformer-Spike 0.269 0.145 1.052 0.620 0.716 0.429 0.813 0.481 0.713 0.419
SIV (ours) 0.234 0.155 0.880 0.573 0.598 0.404 0.677 0.438 0.598 0.392
Table 3: EPE ($\downarrow$) results on PSSD dataset (Problem 3). The best results are marked in bold.
Ablation Study
Ablation of modules. We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 4. All the values are the arithmetic mean of the scenes. The best results are marked in bold. These results underscore the effectiveness of the DPHT, GE, MSVR and $L_{\text{grad}}$. In Table 4, the comparison between experiments (A) and (B) demonstrates the effectiveness of DPHT. Both the comparison between (B) and (C) and that between (D) and (E) demonstrate the effectiveness of the Graph Encoder (GE). Both the comparison between (B) and (D) and that between (C) and (E) highlight the contribution of MSVR.
Ablation of $\beta$. Table 5 summarizes the ablation study on the weighting coefficient $\beta$ of the proposed loss function. The results indicate that setting $\beta=0.3$ yields the best performance for the current model and dataset. Additional ablation studies are provided in the supplementary material.
# Settings Problem 1 Problem 2 Problem 3
DPHT GE MSVR $\Delta t$=21 $\Delta t$=21 $\Delta t$=21
(A) $\times$ $\times$ $\times$ 0.805 1.450 0.776
(B) $\checkmark$ $\times$ $\times$ 0.701 1.299 0.659
(C) $\checkmark$ $\checkmark$ $\times$ 0.625 1.167 0.616
(D) $\checkmark$ $\times$ $\checkmark$ 0.644 1.268 0.635
(E) $\checkmark$ $\checkmark$ $\checkmark$ 0.607 1.144 0.598
Table 4: Ablation study of proposed modules, showing results for EPE ($\downarrow$). $\checkmark$ indicates module presence, $\times$ indicates absence. Setting (G) is our final model.
# $\beta$ Problem 1 Problem 2 Problem 3
(A) 0.0 0.620 1.169 0.615
(B) 0.1 0.617 1.163 0.605
(C) 0.2 0.615 1.157 0.603
(D) 0.3 0.607 1.144 0.598
(E) 0.4 0.613 1.146 0.610
(F) 0.5 0.619 1.141 0.617
Table 5: Ablation study of $\beta$, showing results for EPE (↓).
Conclusion
We explore a novel approach to fluid motion estimation using spike cameras and propose a specialized network architecture, termed Spike Imaging Velocimetry (SIV), designed to solve PIV tasks while adapting to the scenario-specific characteristics of fluid scenes. We propose a Detail-Preserving Hierarchical Transform module to preserve particle signal information during the pre-processing of spike streams. We design a Graph Encoder to extract contextual features from highly unstructured and turbulent flow fields. We develop a Multi-scale Velocity Refinement module to refine the velocity field. We also introduce a challenging and realistic PIV dataset, named Particle Scenes with Spike and Displacement (PSSD). Experimental results demonstrate that our method achieves state-of-the-art performance on the PSSD dataset and accurately estimates velocities in small-scale yet critical flow structures such as shear layers and vortices. In summary, this study demonstrates the potential of spike cameras in fluid velocity measurement and offers new perspectives and directions for PIV network design.